what is the geometric mean between4 and 9
1
Expert's answer
2012-10-23T11:16:29-0400
By definition the geometric mean between numbers A and B is the square root of
their product:
geom_mean(A,B) = sqrt(A*B).

Therefore

geom_mean(4,9) = sqrt(4*9) = sqrt(36) = 6